OpenStudy (anonymous):

Solve -17<4x-6<-1 the solution is{ x|?

OpenStudy (anonymous):

-11/4<x<5/4

jimthompson5910 (jim_thompson5910):

-17<4x-6<-1 -17+6<4x<-1+6 -11<4x<5 -11/4<x<5/4

OpenStudy (anonymous):

is tht a fraction

OpenStudy (anonymous):

yup-11/4 and 5/4 are fractions

jimthompson5910 (jim_thompson5910):

solution really looks like \[\Large \left\{x|-\frac{11}{4}<x<\frac{5}{4}\right\}\]

OpenStudy (anonymous):

ok tht what i was thinking

OpenStudy (anonymous):

ty jim

jimthompson5910 (jim_thompson5910):

yw

OpenStudy (anonymous):

the < have a greater than or equal to signs

jimthompson5910 (jim_thompson5910):

you mean there's a line under the <?

OpenStudy (anonymous):

yes Jim

jimthompson5910 (jim_thompson5910):

so then it's \[\Large \left\{x|-\frac{11}{4}\le x \le \frac{5}{4}\right\}\]

OpenStudy (anonymous):

correct

OpenStudy (anonymous):

ty